OpenStudy (anonymous):

Find the vertex of the graph of the function. f(x) = 2x2 - 8x + 6

OpenStudy (anonymous):

( 2, -2) ( 3, -1) ( -1, 3) ( -2, 2)

OpenStudy (anonymous):

i got ( 3, -1) am i right

OpenStudy (anonymous):

-StudyShark- plez help

zepdrix (zepdrix):

Hey c:

zepdrix (zepdrix):

\[\Large\rm f(x)=2x^2-8x+6\]Need a 1 coefficient on the square term. So factoring out a 2 from each of the x terms gives us:\[\Large\rm f(x)=2(\color{orangered}{x^2-4x})+6\]We need to complete the square on this orange part:\[\Large\rm \left(\frac{b}{2}\right)^2=\left(\frac{-4}{2}\right)^2=(-2)^2=4\]This is the value that completes the square for us. So we'll add 4 to complete our square. But we'll also subtract 4 to keep things balanced.\[\Large\rm f(x)=2(\color{orangered}{x^2-4x+4}-4)+6\]Bringing the -4 out of the brackets, apply the 2 to it,\[\Large\rm f(x)=2(\color{orangered}{x^2-4x+4})-8+6\]Then our perfect square will simplify down to: \(\Large\rm \left(x+\frac{b}{2}\right)^2\)

zepdrix (zepdrix):

\[\Large\rm f(x)=2(x-2)^2-8+6\]

zepdrix (zepdrix):

So it looks like we end up with:\[\Large\rm f(x)=2(x-2)^2-2\]

zepdrix (zepdrix):

Any of those steps look too confusing? :o

OpenStudy (anonymous):

ohh so ir would be (2,-2)

OpenStudy (anonymous):

it*

zepdrix (zepdrix):

Yes.

OpenStudy (anonymous):

can i as another question

zepdrix (zepdrix):

sure.

OpenStudy (anonymous):

State how many imaginary and real zeros the function has. f(x) = x4 - 8x3 + 17x2 - 8x +16 0 imaginary; 4 real 2 imaginary; 2 real 3 imaginary; 1 real 4 imaginary; 0 real i think the answer is 0 imaginary; 4 real

OpenStudy (anonymous):

but i wanted to be sure

zepdrix (zepdrix):

Ummm this involves using De'Cartes Rule of Signs I think. You plug in -x for your x's and count the sign changes I think? Mmm I better go look it up a sec.

zepdrix (zepdrix):

Hmm I'm not sure how they want you to figure this problem out.... De'Cartes Rule of Signs tells us how many real roots we `potentially` could have. Counting the sign changes: We can have `four`, `two`, or `zero` positive real roots. You'll have no negative roots when you apply the second part of the theorem. But then to determine further... I guess you have to apply the Rational Root Theorem maybe? Check all the factors of 16 to see which one will be a root of the polynomial, and then go through the long process of long division. Maybe there is a better way, ugh... But it turns out that it's `2 real`, `2 complex`. How did you come to the conclusion of `4 real`, `0 complex`?

OpenStudy (anonymous):

thanks...i think my answer is right i just wanted to make sure

OpenStudy (anonymous):

thank you for all your help

zepdrix (zepdrix):

But it's not right :o Did you read what I said?